a.. Differentiate the function 𝑦 = (3π‘₯^2+2) ^2√6π‘₯+2 / π‘₯ ^3+1


b. Let P and Q be points on the curve 𝑦 = π‘₯ 2 βˆ’ 2π‘₯ while x = 2 and x = 2 + h respectively. Express the gradient of pq in terms of h and hence find the gradient of the curve 𝑦 = π‘₯ 2 βˆ’ 2π‘₯ at x= 2


c. Find the gradient of the curve 𝑦 = 1/ π‘₯βˆ’1 at the point (2,1)


A straight line 𝑦 = βˆ’π‘₯ + 4 cuts the parabola with equation 𝑦 = 16 βˆ’ π‘₯ 2 at the points A and B.


a) Find the coordinates of A and B


b) Calculate the distance between the points A and B


c) Find the equation of the tangents at A and B, and hence determine where the tangents meet.


d) The line Β΅ is perpendicular to the line A at the point A and B meet. Give its equation


1
Expert's answer
2021-04-15T06:50:15-0400

1.

a.


"y'=((3x^2+2)^2\sqrt{6}x+\dfrac{2}{x^3+1})'"

"=12\sqrt {6}x^2(3x^2+2)+\sqrt{6}(3x^2+2)^2-\dfrac{6x^2}{(x^3+1)^2}"

b.


"grad=\dfrac{(2+h)^2-2(2+h)-2^2+2(2)}{2+h-2}"

"=\dfrac{4+4h+h^2-4-2h-4+4}{h}=2+h"


"\lim\limits_{h\to 0}\dfrac{(2+h)^2-2(2+h)-2^2+2(2)}{2+h-2}"

"=\lim\limits_{h\to 0}(2+h)=2+0=2"

Gradient of the curve "y=x^2-2x" at "x=2" is 2.


c.


"y'=(\dfrac{1}{x-1})'=-\dfrac{1}{(x-1)^2}"

"grad|_{(2, 1)}=-\dfrac{1}{(2-1)^2}=-1"


2.

a) Find the coordinates of A and B


"-x+4=16-x^2"

"x^2-x-12=0"

"(x+3)(x-4)=0"

"x_1=-3, y_1=-(-3)+4=7"


"x_2=4, y_2=-4+4=0"


"A(-3, 7), B(4, 0)"


b)


"d_{AB}=\sqrt{(4-(-3))^2+(0-7)^2}=2\sqrt{7}"

c)


"(16-x^2)'=-2x"

"A(-3, 7)"


"y'(-3)=-2(-3)=6"

"y-7=6(x-(-3))"

The equation of the tangent line

"y=6x+25"

"B(4, 0)"


"y'(4)=-2(4)=-8"

"y-0=-8(x-4)"

The equation of the tangent line

"y=-8x+32"


"6x+25=-8x+32"

"14x=7"

"x=\dfrac{1}{2}"

"y=-8(\dfrac{1}{2})+32=28"

The tangents meet at the point "(\dfrac{1}{2}, 28)"


d)

The line Β΅ is perpendicular to the line A is perpendicular to the line "y=-x+4"


"slope=1"

"y=x+b"

The line Β΅ passes through the point "(\dfrac{1}{2}, 28)"


"28=\dfrac{1}{2}+b=>b=\dfrac{55}{2}"

The equation of the line Β΅ is "y=x+\dfrac{55}{2}."